The supplement of an angle is 60 degrees less than twice the supplement of the complement of the angle.
Answer by nerdybill(6948)

The supplement of an angle is 60 degrees less than twice the supplement of the complement of the angle.
Let x = the angle
then
90-x = complement of angle
180-x = supplement of angle
180-(90-x) = 90+x = supplement of complement of angle
.
180-x = 2(90+x) - 60
180-x = 180+2x -60
-x = 2x -60
-x+60 = 2x
60 = 3x
20 degrees = x